Write the standard equation for the circle. Then state the coordinates of its center and give its radius.
X^2+y^2=4y=12

Answer by lwsshak3(11628)

Write the standard equation for the circle. Then state the coordinates of its center and give its radius. X^2+y^2=4y=12
**
Assume the "=" sign before the 4y should be "-"
Standard equation of a circle of the form: (x-h)^2+(y-k)^2=r^2, (h,k) being the (x,y) coordinates of the center, r=radius
x^2+y^2-4y=12
completing the square
x^2+(y^2-4y+4)=12+4=16
x^2+(y-4)^2=16
For given equation of a circle:
center:(0,4)
radius=√16=4